OpenStudy (anonymous):

I need help! What are the solutions of the system? y= x^2 - 2x - 2 y= 4x + 5

OpenStudy (anonymous):

a. (-1,1) and (-7,-23) b. (-1,1) and (7,23) c. (-1,33) and (7, 1) d. no solution

OpenStudy (turingtest):

substitute the expression fo y from the first equation into the second, giving\[4x+5=x^2-2x-2\]which can be put into the form\[ax^2+bx+c=0\]and solved by factoring, completing the square, or the quadratic formula.

OpenStudy (turingtest):

for*

OpenStudy (anonymous):

I am confused....

OpenStudy (turingtest):

if both \(a=c\) and \(b=c\) then \(a=b\) does that make sense?

OpenStudy (anonymous):

not really... sorry I am so bad at algebra

OpenStudy (turingtest):

well this is not an easy problem if you are "bad" at algebra (don't jump to conclusions, I felt that way once too) do you know the quadratic formula?

OpenStudy (anonymous):

yes... kinda

OpenStudy (anonymous):

ax2 + bx + c", right?

OpenStudy (turingtest):

the quadratic formula is that for equations of the form\[ax^2+bx+c=0\]we have\[x={-b\pm\sqrt{b^2-4ac}\over2a}\]

OpenStudy (turingtest):

now we want to make a substitutionsince both expressions with x are equal to y we can substitute